OpenStudy (anonymous):

Solve the equation, using the domain of x = {-1, 0, 2}. 2x + 3y = 7 A. {(-1, 3), (0, 7/3), (2, 1)} B. {(3, -1), (7/3, 0), (1, 2)} C. {(5, -1), (7/2, 0), (2, 1)} D. {(-1, -3), (0, 7/3), (2, 1)}

OpenStudy (the_fizicx99):

Yup ok so plug in those x values into the equation 2x + 3y = 7 2(-1) + 3y = 7 -2 + 3y = 7 +2 +2 3y = 9 /3 /3 y = 3 So, (-1,3) for the x value of -1

OpenStudy (the_fizicx99):

Continue plug in those values and you'll narrow it down :p